OpenStudy (anonymous):

-4c- 7c-(-8+c) simplify

OpenStudy (anonymous):

i got -3

OpenStudy (anonymous):

-11c+8-c=-12c+8

OpenStudy (anonymous):

how 3?

OpenStudy (anonymous):

i add the 4 and 7 together

OpenStudy (anonymous):

and then - 8

OpenStudy (anonymous):

-4c-7c=-11c, ok?

OpenStudy (anonymous):

-11c-(-8+c) =-11c+8-c =-12c+8

OpenStudy (anonymous):

ok